OpenStudy (dmezzullo):

Explain how to graph 2x + y = 14 using the x- and y-intercepts method. Be sure to use complete sentences and use the x- and y-intercepts method only.

OpenStudy (saifoo.khan):

Two sentences: 1) When x-intercepts, make y=0 2) When y-intercepts, make x=0

OpenStudy (dmezzullo):

Thats it?

OpenStudy (saifoo.khan):

Yes. Now rewrite this in your own words with your particular problem.

OpenStudy (dmezzullo):

Wut u mean my particular problem?

OpenStudy (saifoo.khan):

2x + y = 14

OpenStudy (dmezzullo):

Oh ok Ik how i would do da doe

OpenStudy (dmezzullo):

idk*

OpenStudy (saifoo.khan):

2x + y = 14 1) When x-intercepts, make y=0 2x + (0) = 14 x = 14/2 x = 7 So x-int is 7. 2) When y-intercepts, make x=0 now find the y-intercept.

OpenStudy (anonymous):

Anyone here have the formula for the area of a circle? i forgot it

OpenStudy (dmezzullo):

I=prt?

OpenStudy (saifoo.khan):

@Isaacclarke2 Area = pi*r^2

OpenStudy (anonymous):

^Thx Khan